OpenStudy (anonymous):

For g(x)=4-5x, determine the input for "x" when the output of g(x) is 3/5. Explain.

OpenStudy (anonymous):

4-5x = 3/5 -5x = 3/5 - 4 -5x = -22/5 5x = 22/5 x = 22/25